Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by rbansal » Sun Apr 10, 2011 5:06 pm
GMAT A bag holds 2 red marbles and 3 green marbles. If you removed two randomly selected marbles from the bag, what is the probability that both would be red?

ANSWER CHOICES

A 1/10
B 1/5
C 3/10
D 2/5
E 1/2


I eliminated D & E right away but the answer is A and I dont understand how they got there. Can you give me a detailed step by step explanation.
Join the discussion
Source: — Problem Solving |

by Stuart@KaplanGMAT » Sun Apr 10, 2011 6:12 pm
rbansal wrote:GMAT A bag holds 2 red marbles and 3 green marbles. If you removed two randomly selected marbles from the bag, what is the probability that both would be red?

ANSWER CHOICES

A 1/10
B 1/5
C 3/10
D 2/5
E 1/2


I eliminated D & E right away but the answer is A and I dont understand how they got there. Can you give me a detailed step by step explanation.
Hi,

when you select multiple items simultaneously, you can treat the question as though you select them one at a time, without replacement.

So, our question becomes:

If there are 2 red marbles and 3 green marbles in a bag, and you select a marble, then select a second marble without replacing the first, what's the probability that both marbles selected were red?

Now it's just simple probability:

prob = (# of desired outcomes)/(total # of possibilities)

Prob 1st marble red = 2/5
Prob 2nd marble red = 1/4

(without replacement, so we adjust the numbers as we go).

Since we want the probability that BOTH are red, we MULTIPLY each independent probability:

(2/5)(1/4) = 2/20 = 1/10

Presto!
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by vineeshp » Sun Apr 10, 2011 6:13 pm
The answer has to be 2C2/5C2.

1 / 10.

A.

Detailed explanation

No. of possible outcomes : From 5 marbles (2 red and 3 green), you can choose 2. That is 5C2 which is equal to 10.

No of desired outcomes: Both have to be red. From 2, select 2. There is only one way of doing this. That is to select both. (2C2)

So Probablity = Desired outcomes / All outcomes.

1 / 10.
Vineesh,
Just telling you what I know and think. I am not the expert. :)
Join the discussion

by rbansal » Sun Apr 10, 2011 6:31 pm
Thank you so much this makes much more sense.

RB
Join the discussion

by jacob.t.jennings » Tue May 24, 2011 3:42 pm
@ Stuart,

On probability questions, I realized that you sometimes use the "one-minus" method and sometimes you don't. For example, on this question, you did not. I'm guess it was because it's not as complex as some of the other questions? So, do you have a recommendation on what type of probability question to use the "one-minus" method and when not to? I would appreciate your response.
Join the discussion

by Stuart@KaplanGMAT » Tue May 24, 2011 6:36 pm
jacob.t.jennings wrote:@ Stuart,

On probability questions, I realized that you sometimes use the "one-minus" method and sometimes you don't. For example, on this question, you did not. I'm guess it was because it's not as complex as some of the other questions? So, do you have a recommendation on what type of probability question to use the "one-minus" method and when not to? I would appreciate your response.
Great question!

The "one-minus" method is most useful on multiple scenario questions in which there are more scenarios that we want than we don't want.

The question will often have "at least" or "at most" in it, indicating that there are multiple possibilities that you want and don't want.
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by jacob.t.jennings » Wed May 25, 2011 11:04 am
Thanks a lot Stuart!
Join the discussion

by Vishnu88 » Wed May 25, 2011 11:10 am
Stuart, would you be kind enough to illustrate a probability problem with multiple scenarios. Just to get an idea of the one-minus method?
Join the discussion

by Stuart@KaplanGMAT » Wed May 25, 2011 10:49 pm
Vishnu88 wrote:Stuart, would you be kind enough to illustrate a probability problem with multiple scenarios. Just to get an idea of the one-minus method?
Sure! Let's use a very common GMAT probability question type: coin flips.
A fair coin is flipped 5 times. What's the probability that at least 1 of the 5 results is heads?
Whenever you see "at least" or "at most" in a question, you're dealing with multiple scenarios. Here, for example, we want "at least 1 head" out of 5 flips, so acceptable results are:

1 head and 4 tails;
2 heads and 3 tails;
3 heads and 2 tails;
4 heads and 1 tail; and
5 heads and 0 tails.

Since we want 1H4T OR 2H3T OR 3H2T OR 4H1T OR 5H01, one way to solve would be to figure out the probability of each scenario and ADD them together.

(Here's a very important rule to remember for probability and counting questions: to calculate MULTIPLE scenarios, MULTIPLY the individual ones; to calculate ALTERNATIVE scenarios, ADD the individual ones.)

However, if we're GMAT experts we notice that the only scenario that we DO NOT want is 0 heads and 5 tails. Since the sum of all probabilities is 1, we can use the "one minus" approach and set up the following equation:

Probability(what we want) = 1 - Probability(what we don't want)

Prob(at least 1 head) = 1 - Prob(0 heads)

0 heads is the same as 5 tails, and the probability of 5 tails is simply:

(1/2)(1/2)(1/2)(1/2)(1/2) = 1/32

So:

Prob(at least 1 head) = 1 - 1/32 = 31/32
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by Vishnu88 » Wed May 25, 2011 10:56 pm
Stuart, thank you so much for that.

Couldn't have asked for anything better.

Cheers!
Join the discussion